# OpenPros: A Large-Scale Dataset for Limited View Prostate Ultrasound Computed Tomography

### v1

@@ 4 OpenPros Benchmarks
Table 4: Quantitative results of benchmarking methods on the prostate USCT dataset.
[…]
InversionNet 0.0089 0.0433 0.9845 0.9798
ViT-Inversion 0.0123 0.0566 0.9774 0.9728
OpenPros lets us systematically investigate three core questions under limited-angle prostate USCT: (1) Inference efficiency, (2) Reconstruction accuracy, and (3) OOD generalization.
Methods & Metrics
We compare two physics-based baselines-Delay-and-Sum Beamforming and multi-stage physics-based USCT inversion-against two data-driven models: InversionNet, and ViT-Inversion. Performance is measured with four metrics: mean absolute error (MAE) and root-mean-square error (RMSE) for numerical fidelity, plus SSIM and Pearson correlation (PCC) for perceptual and structural similarity. All models are trained and evaluated under identical settings on NVIDIA H100 GPUs.
@@ 4.1 Benchmark Methods for Prostate USCT
Figure 4: Benchmark results for limited-angle prostate USCT. Each column shows a different inversion method on the same phantom: (col 1) ground-truth SOS map; (col 2) Delay-and-Sum beamforming; (col 3) physics-based USCT; (col 4) InversionNet; (col 5) ViT-Inversion. Rows correspond to three representative prostate slices illustrating challenging (top), moderate (middle), and simple (bottom) anatomical scenarios. Zoom-in figures of the prostate region (orange squares) are shown in Figure 5.
@@
Physics-based USCT is perform in a three-stage multi-frequency framework. Starting from a smoothed initial SOS model, we first invert low-frequency data, then mid-frequency data, and finally full-band data. At each stage, synthetic waveforms are generated via our forward operator and compared to observed data; the SOS model is updated by minimizing the waveform misfit.
InversionNet wu2019inversionnet proposed a fully-convolutional network to model the seismic inversion process. With the encoder and the decoder, the network was trained in a supervised scheme by taking 2D (time $\times$ # of receivers) seismic data from multiple sources as the input and predicting 2D (depth $\times$ length) velocity maps as the output.
ViT-Inversion: A Vision Transformer that partitions the 4D waveform tensor $[B,S,T,R]$ into spatio-temporal patches, embeds them into tokens, and applies multi-head self-attention to capture long-range wave interactions. A lightweight upsampling CNN refines the patch-wise outputs into full-resolution SOS maps.
(Training details in App. E.)
@@ 4.2 Result Analysis
Table 4 reports mean absolute error (MAE), root mean squared error (RMSE), structural similarity index (SSIM), and Pearson correlation coefficient (PCC) for our two learned baselines. Traditional physics-based USCT (not shown) achieves an RMSE around 0.16 and SSIM $\sim 0.90$, roughly halving the error of simple beamforming but still leaving significant room for improvement. In contrast, InversionNet reduces RMSE to 0.0433-nearly a fourfold improvement over physics-based inversion-and attains the highest MAE (0.0089), SSIM (0.9845), and PCC (0.9798). ViT-Inversion also outperforms the physics-based approach, cutting RMSE in half (0.0566) and achieving an SSIM of 0.9774.
@@ 5 Ablation Study
To assess how well our models generalize to truly unseen anatomies, we conducted three out-of-distribution tests using data splits that reflect realistic clinical scenarios: (1) Patient-Level Generalization: Train on patients 3_01, 3_02, 3_03, test on an entirely unseen patient 3_04. (2) Leave-One-Prostate-Out: Train on 60 prostates from all four patients except two held-out prostate (i.e. 2022-06-06 & 2022-06-09), test on those withheld prostates. (3) Combined Generalization: Train on 60 prostates drawn only from patients 3_013_03, test on both remaining prostates of patient 3_04.
Table 5 summarizes MAE, RMSE, SSIM, and PCC for InversionNet and ViT-Inversion under each scenario. In the patient-level generalization test, all methods suffer substantial performance degradation. InversionNet’s MAE increases from 0.0089 (within-distribution) to 0.0653 and its RMSE from 0.0433 to 0.2465, while SSIM falls to 0.9111. ViT-Inversion experiences a similar drop, with MAE rising to 0.0802, RMSE to 0.2898, and SSIM to 0.9055. This indicates that entirely novel patient anatomies pose a significant challenge for learned inversion models.
By contrast, the leave-one-prostate-out test shows only minimal accuracy loss. InversionNet maintains MAE=0.0070, RMSE=0.0291, and SSIM=0.9886, nearly matching in-distribution performance. ViT-Inversion also remains strong with MAE=0.0100, RMSE=0.0424, and SSIM=0.9823. These results suggest that intra-patient anatomical variability is comparatively easier for the networks to handle.
The most demanding combined OOD scenario amplifies the accuracy drop. InversionNet’s SSIM settles at 0.9137 (MAE=0.0624, RMSE=0.2370), and ViT-Inversion achieves SSIM=0.8972 (MAE=0.0862, RMSE=0.3045). Although InversionNet retains slightly higher fidelity under these extreme conditions, neither method reaches within-distribution levels, underscoring that truly patient-agnostic reconstruction remains an open challenge.
@@
Patient-level InversionNet 0.0653 0.2465 0.9111 0.6574
ViT-Inversion 0.0802 0.2898 0.9055 0.6098
Leave-one-prostate InversionNet 0.0070 0.0291 0.9886 0.9891
ViT-Inversion 0.0100 0.0424 0.9823 0.9852
Combined OOD InversionNet 0.0624 0.2370 0.9137 0.6753
ViT-Inversion 0.0862 0.3045 0.8972 0.5774
Overall, these ablations demonstrate that while machine learning based methods significantly outperform physics-based USCT under matched conditions, their ability to generalize to entirely unseen anatomies is still limited.
@@ 6 Discussion
However, the dataset currently includes a limited number of patient anatomies, potentially underrepresenting certain anatomical variations. Additionally, we simulate only SOS distributions, omitting other critical acoustic parameters like attenuation and density. Our 2D simulations do not account for three-dimensional propagation and out-of-plane scattering, simplifying some real-world conditions.
@@ 7 Conclusion
  • Section A highlights the public resources to support the reproducibility and describes the licenses of the OpenPros data and released code.
  • Section B describes the detailed steps of how 2D phantoms are sliced from the 3D speed of sound volumes.
  • Section C shows the naming strategy of the OpenPros dataset.
  • Section D describes the benchmark evaluation metrics of OpenPros.
  • Section E shows the baseline models’ training configurations and hyper-parameters.
  • Section F compares the widely used conventional k-Wave simulation method and our OpenPros open-sourced simulation methods.
Appendix A OpenPros Public Resources and Licenses
First and foremost, the reproducibility of OpenPros benchmarks is guaranteed by a number of public resources, listed below. Remarkably, we have a group (link available below) where any related discussion is welcome. Our team also promises to maintain the platform and support further developments based on the community feedback.
The codes are released on Github under OSS license and BSD-3 license. We also attach the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License to the data.
Appendix B Details of 2D Phantom Extraction
@@
Altogether, this procedure produces 280 000 unique 2D SOS phantoms with matching ultrasound data under limited-angle conditions. An illustration of the slicing geometry is shown in Figure 6.
Appendix C Naming of OpenPros
@@ Appendix E Training Procedure
Appendix F Conventional k-Wave vs. Our Ultrasound Simulation Algorithms
Ultrasound imaging methods can be broadly categorized by their simulation paradigms. Conventional TRUS simulations, commonly used for B-mode imaging, rely on signal processing pipelines such as delay-and-sum (DAS) beamforming applied to simulated echoes. These simulations are often implemented using toolboxes like MATLAB k-Wave [24], which model acoustic wave propagation using pseudospectral methods and reconstruct images from envelope-detected signals. While fast and widely adopted in the clinical ultrasound community, this approach simplifies underlying physics and often introduces artifacts due to assumptions like homogeneous backgrounds, limited diffraction modeling, or approximate transducer responses.

### v2

@@ 4 OpenPros Benchmarks
OpenPros enables systematic investigation of three core questions in limited-angle prostate USCT: (1) inference efficiency, (2) reconstruction accuracy, and (3) out-of-distribution (OOD) generalization. We compare two physics-based baselines, Delay-and-Sum beamforming and multi-stage USCT inversion, against two data-driven models: CNN-based InversionNet (Wu and Lin, 2019) and a Vision Transformer (ViT)-based variant (Dosovitskiy et al., 2020), referred to here as ViT-Inversion. Performance is evaluated using four metrics: mean absolute error (MAE) and root-mean-square error (RMSE) for numerical fidelity, and structural similarity index (SSIM) and Pearson correlation coefficient (PCC) for perceptual and structural alignment. All models are trained and evaluated under identical settings on NVIDIA H100 GPUs.
@@ 4.1 Benchmark Methods for Prostate USCT
Figure 4: Benchmark results for limited‐angle prostate USCT. Each column shows a different inversion method on the same phantom: (col 1) ground‐truth SOS map; (col 2) Delay‐and‐Sum beamforming; (col 3) physics‐based USCT; (col 4) InversionNet; (col 5) ViT‐Inversion. Rows correspond to three representative prostate slices illustrating challenging (top), moderate (middle), and simple (bottom) anatomical scenarios. Zoom-in figures of the prostate region (orange squares) are shown in Figure 5.
@@
Physics-based USCT is performed in a three‐stage multi-frequency framework. Starting from a smoothed initial SOS model, we first invert low‐frequency data, then mid‐frequency data, and finally full‐band data. At each stage, synthetic waveforms are generated via our forward operator and compared to observed data; the SOS model is updated by minimizing the waveform misfit.
InversionNet (Wu and Lin, 2019) proposed a fully-convolutional network to model the seismic inversion process. With the encoder and the decoder, the network was trained in a supervised scheme by taking 2D (time $\times$ # of receivers) seismic data from multiple sources as the input and predicting 2D (depth $\times$ length) velocity maps as the output.
ViT-Inversion: A Vision Transformer (Dosovitskiy et al., 2020) that partitions the 3D waveform tensor $[S,T,R]$ into spatio‐temporal patches, embeds them into tokens, and applies multi‐head self‐attention to capture long‐range wave interactions. A lightweight upsampling CNN refines the patch‐wise outputs into full‐resolution SOS maps.
The train/test splits for the in-distribution (ID) experiments in the section: we randomly partition all available 2D samples at the slice level into training, and test sets (90% / 10%), without enforcing patient exclusivity. As a result, slices from the same 3D anatomy can appear in both train and test sets; this protocol is intentionally designed to measure interpolation performance within the set of anatomies seen during training.
Training configurations and hyperparameters are provided in App. A.4.
@@ 4.2 Result Analysis
Table 4: Quantitative in-distribution results (mean $\pm$ std over all test slices).
[…]
InversionNet $0074\pm 0.0037 0.0$$0247\pm 0.0166$ $0.9955\pm 0.0037$ $0.9851\pm 0.0262$
ViT-Inversion $\mathbf{0.0067\pm 0.0057}$ $\mathbf{0.0205\pm 0.0172}$ $\mathbf{0.9967\pm 0.0035}$ $\mathbf{0.9893\pm 0.0219}$
[…]
Qualitative Observations
@@ 5 Ablation Study
Generalization Tests
To assess how well our models generalize to truly unseen anatomies, we conducted three out‐of‐distribution tests using data splits that reflect realistic clinical scenarios: (1) Patient-Level Generalization: Train on patients 3_01, 3_02, 3_03, test on an entirely unseen patient 3_04. (2) Leave-One-Prostate-Out: Train on 60 prostates from all four patients except two held‐out prostates (i.e., 2022-06-06 & 2022-06-09); test on those withheld prostates. (3) Combined Generalization: Train on 60 prostates drawn only from patients 3_013_03; test on both remaining prostates of patient 3_04.
[…]
@@
Patient-level InversionNet 0.0322 0.1010 0.9399 0.8271
ViT-Inversion 0.0276 0.0890 0.9496 0.8689
Leave-one-prostate InversionNet 0.0069 0.0273 0.9899 0.9894
ViT-Inversion 0.0061 0.0210 0.9934 0.9937
Combined OOD InversionNet 0.0323 0.1017 0.9408 0.8251
ViT-Inversion 0.0280 0.0916 0.9482 0.8663
By contrast, the leave-one-prostate-out split shows minimal and, in some cases, negligible degradation compared to ID. ViT-Inversion achieves slight improvements over InversionNet across all metrics. These results show that intra-patient anatomical variability poses little difficulty for the models, thus generalization across unseen prostates is comparatively straightforward.
The combined OOD split mirrors the patient-level challenge: both models exhibit substantial error increases, with ViT-Inversion consistently outperforming InversionNet across metrics. Overall, while intra-patient generalization is well handled, robust patient-agnostic reconstruction under limited-angle conditions remains a significant open challenge.
Robustness to Input Noise
We also assess robustness to measurement corruption by adding zero-mean Gaussian noise to the test waveforms while keeping all models trained on clean data only; see App. A.6 for full protocol and tables. We sweep $\sigma\in\{0.01,0.02,0.05\}$ (roughly $26/23/19\,\mathrm{dB}$ PSNR). Performance decreases monotonically with noise, and ViT-Inversion is consistently more resilient (SSIM $0.987\!\rightarrow\!0.935$) than InversionNet (SSIM $0.944\!\rightarrow\!0.825$). Details and additional metrics (MAE/MSE/RMSE/PCC) are reported in App. A.6.
@@ 6 Discussion
However, the dataset currently includes a limited number of patient anatomies, potentially underrepresenting certain anatomical variations. Additionally, we simulate only SOS distributions, omitting other critical acoustic parameters like attenuation and density. Our 2D simulations do not account for three-dimensional propagation and out-of-plane scattering, simplifying some real-world conditions. This 2D, SOS-only setting is a clear simplification compared to a full 3D clinical USCT system, but it preserves the key ill-posedness of limited-angle prostate USCT while remaining computationally tractable. It enables us to generate approximately 280k paired waveform-SOS samples and to systematically compare reconstruction methods at scale. We therefore position OpenPros as a fast-prototyping benchmark for waveform-to-SOS reconstruction, with the expectation that successful ideas will later be re-evaluated in more complete 3D, multi-parameter models.
Beyond the baseline CNN and ViT models evaluated in this paper, OpenPros is designed to serve as a testbed for a broader family of reconstruction approaches, including UNet-like encoder–decoder architectures, diffusion and other generative models, and neural operator methods that directly learn PDE solution operators (e.g., (Dai et al., 2023)). The large number of paired waveform–SOS examples and the availability of both ID and OOD evaluation protocols make it possible to systematically study how these different model classes trade off reconstruction fidelity, generalization across anatomies, and computational cost in realistic limited-angle USCT settings.
@@ Appendix A Appendix
  • Section A.1 describes the detailed steps of how 2D phantoms are sliced from the 3D speed of sound volumes.
  • Section A.2 shows the naming strategy of the OpenPros dataset.
  • Section A.3 describes the benchmark evaluation metrics of OpenPros.
  • Section A.4 shows the baseline models’ training configurations and hyper-parameters.
  • Section A.5 compares the widely used conventional k-Wave simulation method and our OpenPros open-sourced simulation methods.
A.1 Details of 2D Phantom Extraction
@@
Altogether, this procedure produces 280,000 unique 2D SOS phantoms with matching ultrasound data under limited-angle conditions. An illustration of the slicing geometry is shown in Figure 6.
A.2 Naming of OpenPros
@@ A.4 Training Procedure
A.5 Conventional k-Wave vs. Our Ultrasound Simulation Algorithms
Ultrasound imaging methods can be broadly categorized by their simulation paradigms. Conventional TRUS simulations, commonly used for B-mode imaging, rely on signal processing pipelines such as delay-and-sum (DAS) beamforming applied to simulated echoes. These simulations are often implemented using toolboxes like MATLAB k-Wave (Treeby and Cox, 2010), which model acoustic wave propagation using pseudospectral methods and reconstruct images from envelope-detected signals. While fast and widely adopted in the clinical ultrasound community, this approach simplifies underlying physics and often introduces artifacts due to assumptions like homogeneous backgrounds, limited diffraction modeling, or approximate transducer responses.
@@
A.6 Noise Robustness under Additive Measurement Perturbations
Setup
To assess robustness to measurement corruption, we perturb the test waveforms with zero-mean i.i.d. Gaussian noise and evaluate the pretrained models without any fine-tuning (all models were trained only on clean data):
$\tilde{\mathbf{p}}\;=\;\mathbf{p}\;+\;\bm{\epsilon},\qquad\bm{\epsilon}\sim\mathcal{N}\!\left(\mathbf{0},\,\sigma^{2}\mathbf{I}\right).$
Noise is applied to the input tensor after the same normalization used during training. We sweep $\sigma\in\{0,\,0.01,\,0.02,\,0.05\}$; for reference, these levels correspond to input PSNRs of approximately 26dB26\ $\mathrm${dB}, 23dB23\ $\mathrm${dB}, and 19dB19\ $\mathrm${dB}, respectively. Reconstructions are compared to the clean ground-truth SOS using MAE, MSE, RMSE, SSIM, and PCC.
Results
Tables 8 and 9 summarize performance for ViT-Inversion and InversionNet. Both models degrade monotonically as noise increases, but ViT-Inversion is substantially more resilient across all metrics. At $\sigma=0.01$ ($\approx 26\ \mathrm{dB}$), ViT-Inversion maintains SSIM $\approx 0.987$ and PCC $\approx 0.982$, while InversionNet drops to SSIM $0.944$ and PCC $0.837$. At the highest noise level ($\sigma=0.05$, $\approx 19\ \mathrm{dB}$), ViT-Inversion still preserves moderate structural fidelity (SSIM $0.935$, PCC $0.796$), whereas InversionNet falls to SSIM $0.825$ and PCC $0.388$. These trends suggest that attention-based models better suppress high-frequency perturbations by leveraging longer-range context in the wavefield.
Table 8: Noise robustness of ViT-Inversion. Gaussian noise $\mathcal{N}(0,\sigma^{2})$ added to test inputs; models trained on clean data only.
ViT-Inversion $\sigma=0$ $\sigma=0.01$ $\sigma=0.02$ $\sigma=0.05$
PSNR (dB) 26 23 19
MAE 0.0067 0.0084 0.0137 0.0366
RMSE 0.0268 0.0330 0.0521 0.1149
PCC 0.9893 0.9824 0.9507 0.7965
SSIM 0.9909 0.9872 0.9759 0.9347
Table 9: Noise robustness of InversionNet. Same protocol as Table 8.
InversionNet $\sigma=0$ $\sigma=0.01$ $\sigma=0.02$ $\sigma=0.05$
PSNR (dB) 26 23 19
MAE 0.0074 0.0287 0.0573 0.1228
RMSE 0.0297 0.0964 0.1700 0.3155
PCC 0.9851 0.8372 0.6604 0.3881
SSIM 0.9877 0.9437 0.8998 0.8252
Takeaways and implications
(1) Robustness without noise exposure during training is limited—especially for convolution-only models—highlighting the importance of noise-aware data augmentation and/or denoising front-ends. (2) Attention mechanisms appear to confer greater stability to measurement noise in this task. (3) Practical deployment will likely benefit from simple defenses such as SNR-matched augmentation, temporal filtering of waveforms, and uncertainty-aware inference; we include these as future baselines in subsequent releases.
A.7 View-dependent difficulty: bone vs. no-bone paths
To better understand which acquisition configurations are easier or harder to reconstruct, we stratify the test slices according to whether the dominant ultrasound propagation paths intersect the pelvic bone. Specifically, we group slices into three categories: (i) no bones, where rays from the probe to the prostate region do not intersect bone; (ii) small bones, where bone intersections occur but occupy a relatively small angular extent; and (iii) big bones, where the prostate region is largely shadowed by bone, corresponding to more severe limited-angle geometries.
Table 10 reports the performance of the ViT-Inversion baseline on these three categories in the in-distribution setting. We observe that MAE and RMSE increase as the amount of bone along the propagation paths increases: the no bones group yields the lowest errors, the small bones group shows intermediate errors, and the big bones group exhibits the highest errors. This trend confirms that views with strong bone interference are substantially harder to reconstruct, while bone-free views are comparatively easier. SSIM remains high in all three categories, indicating good overall structural similarity, but amplitude-wise errors are clearly more pronounced in the presence of bone.
Table 10: View-dependent performance of ViT-Inversion on in-distribution test slices, stratified by bone involvement along the propagation paths. Metrics are reported as means over all slices in each group. Lower is better for MAE/MSE/RMSE; higher is better for PCC/SSIM.
Metric No bones (7063) Small bones (6694) Big bones (13603)
MAE 0.0046 0.0059 0.0082
[…]
RMSE 0.0098 0.0213 0.0342
PCC 0.9740 0.9884 0.9976
SSIM 0.9921 0.9908 0.9902
A.8 Usage of Large Language Models (LLMs)
During the preparation of this manuscript, we used a large language model (LLM) to assist with language polishing, structural refinement, and presentation clarity. The LLM provided feedback on phrasing, grammar, and flow, suggested alternatives to reduce redundancy, and generated formatting for tables and equations. All scientific ideas, experiments, analyses, and conclusions were conceived, designed, and carried out entirely by the authors.